mother is 4 times as old as MAy. 5 years ago, she was 7 times as old as may. how old is each 2 years from now.
Answer by ewatrrr(24785)


Hi,
Let x represent May's age. Then the mother's age would be represented by 4x
5 years ago, the mother was 7 times as old as May***
4x - 5 = 7(x-5)
solving for x
4x - 5 = 7x - 35
30 = 3x
10 = x, May's age NOW. the mother is NOW 40 (4*10)
how old is each 2 years from now
May is 12 and the mother is 42
CHECKING OUR ANSWER
5 years ago, the mother was 7 times as old as May***
35 = 7*5